OpenStudy (anonymous):

does anyone know how to use the matrices in system of equations? its really confusing i need help

OpenStudy (anonymous):

x + 2y - z = -1 x - 4y - z = -4 2x - 2y + 3z = 15

OpenStudy (anonymous):

@mathmath333 @cutelilcigarette @Cuanchi could you help